### v3

# and Objects in Video Generation

Abstract
Controlling the movements of dynamic objects and the camera within generated videos is a meaningful yet challenging task. Due to the lack of datasets with comprehensive 6D pose annotations, existing text-to-video methods can not simultaneously control the motions of both camera and objects in 3D-aware manner, resulting in limited controllability over generated contents. To address this issue and facilitate the research in this field, we introduce a Synthetic Dataset for Free-Form Motion Control (SynFMC). The proposed SynFMC dataset includes diverse object and environment categories and covers various motion patterns according to specific rules, simulating common and complex real-world scenarios. The complete 6D pose information facilitates models learning to disentangle the motion effects from objects and the camera in a video. To provide precise 3D-aware motion control, we further propose a method trained on SynFMC, Free-Form Motion Control (FMC). FMC can control the 6D poses of objects and camera independently or simultaneously, producing high-fidelity videos. Moreover, it is compatible with various personalized text-to-image (T2I) models for different content styles. Extensive experiments demonstrate that the proposed FMC outperforms previous methods across multiple scenarios.
Figure 1: The rule-based generation pipeline of videos in the proposed Synthetic Dataset for Free-Form Motion Control (SynFMC). This example generates synthetic video with three objects: (1) The environment asset and it’s matching object assets are selected as the scene elements. (2) The motion types of objects and camera are randomly selected for trajectory generation. (3) The center region shows the resulting 3D animation sequence used for rendering. The rendered video and annotations are demonstrated in the last row.
1 Introduction
Controlling motion dynamics in video generation has received increasing attention [12, 47, 46, 11, 21, 53, 24, 20, 42, 18, 35], as it enables better customization and is crucial in many applications. For example, in filmmaking, directors meticulously choreograph the movements of both actors and the camera. Consequently, precise control over object and camera motions in video offers creative flexibility.
Despite recent progress, challenges remain in motion control of text-to-video (T2V) generation. A key limitation is the lack of high-quality datasets with comprehensive 6D pose annotations. For controlling object movement [47, 23, 44, 54, 59, 25, 27, 4, 48, 22, 34], the motion is primarily annotated as the trajectory in image space [47, 54]. This annotation, however, lacks 3D nature and intertwines the dynamics of both objects and the camera. For example, a rightward trajectory could represent either a stationary camera with a moving object or a static object with a left-moving camera. Recently, 360°360°360  °-Motion synthetic dataset [10] provides 6D poses of objects, but limited with static camera setting and motion diversity. Besides, existing commonly used datasets [51, 61, 57] for learning camera motion mainly focus on scenes with minimal object dynamics. Some human-centric synthetic datasets [46, 3, 52, 55] provide ground truth for both human subjects and camera motions within a global coordinate system, yet exhibit limited motion diversity and category variety.
Another limitation is the absence of methods that can independently or jointly control the 6D poses of both object and camera. For example, methods like Motion-Zero [4] animate objects without 3D-aware control (e.g., orientation) [19, 48, 25], while methods like CameraCtrl [12] exclusively focus on camera motion [16, 49, 1]. MotionCtrl [47] trains separate modules for object and camera control in a two-stage process. However, without access to video data containing complete 6D pose annotations for both elements, it struggles to achieve realistic, synchronized control of objects and cameras within the same scene.
Figure 2: Example videos generated by our method FMC trained on the proposed SynFMC dataset, showing its adaptability with different personalized T2I models [30, 38, 41].
To address these limitations, a dataset with comprehensive 6D pose annotations of objects and camera is desired. However, acquiring such data is challenging and typically requires specialized equipment and expertise. In this work, we introduce a Synthetic dataset for Free-Form Motion Control (SynFMC). Designed with an emphasis on quality and diversity, the dataset includes a rich array of animated object assets and environment assets across various categories. Furthermore, a rule-based generation algorithm is implemented to create trajectories for both objects and the camera, as shown in Fig. 1. This algorithm encompasses basic patterns and simulates challenging cinematographic shots as in Fig. 5. Compared to recent works [10, 46] that can only construct uncontrolled trajectories, our rule-based algorithm supports customized object & camera motions with diverse patterns. To enhance realism, essential attributes of objects, like living environment, types of speed and size, are annotated using a Multimodal Large Language Model (MLLM) [5] and manual labeling, facilitating the generation of plausible trajectories. The SynFMC dataset also provides detailed annotations, including 6D pose information of objects and camera, instance segmentation maps, depth maps, and comprehensive descriptions of content and motion, supporting a wide range of research fields.
To further validate the effectiveness of the proposed SynFMC dataset and support 3D-aware control in T2V generation, we propose a Free-Form Motion Control (FMC) method. FMC mainly includes two components: Camera Motion Controller (CMC) and Object Motion Controller (OMC). Unlike previous methods [54, 47], Our approach trained on SynFMC disentangles global (camera) and local (object) dynamics and manipulates the 6D poses of camera and objects. Furthermore, we adopt Domain LoRA [17] to prevent model from fitting to rendered style in synthetic data. As shown in Fig. 2 and Fig. 6, FMC effectively mitigates the domain gap and adapts to various personalized Text-to-Image (T2I) models, generating high-fidelity results across diverse styles. In addition, FMC provides flexible user interfaces for motion control. Users can input trajectories for objects and camera by simply drawing 3D curves or by specifying motion types for each (as detailed in Sec. 3.3), which are used by the rule-based algorithm to generate their trajectories accordingly.
Table 1: Comparison of the proposed SynFMC with existing datasets. The object/camera motion pattern columns apply only to synthetic datasets. In addition to offering a rich variety of object categories, SynFMC outperforms in motion pattern variety and controllability with comprehensive pose annotations of camera and objects. In our implementation, we only use 26K subset as training data.
Dataset Clips Source Category Object Motion Pattern Camera Motion Pattern Camera Pose Annotation Object Motion Annotation
RealEstate10K [61]   65K Real Real Estate - - Fitting
MVImgNet [57] 220K Real Common - - Fitting
VideoHD [54]   75K Real Common - - Optical Flow
MotionCtrl [47] 240K Real Common - - Optical Flow
HumanVid-Real [46]   20K Real Human - - Fitting 2D Human Pose
BEDLAM [3]   10K Synthetic Human Limited/Uncontrollable Static Ground Truth 3D Human Pose
SynBody [52]   27K Synthetic Human Limited/Uncontrollable Static Ground Truth 3D Human Pose
HumanVid-Syn [46] 100K Synthetic Human Limited/Uncontrollable Diverse/Uncontrollable Ground Truth 3D Human Pose
360°360°360  °-Motion [10]   54K Synthetic Animal & Human Limited/Uncontrollable Static Object Pose
SynFMC (ours)   62K Synthetic Common Diverse/Controllable Diverse/Controllable Ground Truth Object Pose
In summary, our main contributions are as follows:
  • To the best of our knowledge, the SynFMC dataset is the first to provide 6D pose annotations for both camera and objects. Its diverse scenes and complex motion patterns provide models with valuable resources for learning the dynamics of multiple objects and the camera.
  • The FMC method can manipulate the 6D poses of the camera and objects independently or simultaneously, achieving high-quality results across diverse scenes.
  • Extensive experiments demonstrate that FMC, trained on SynFMC dataset, generates videos of superior quality compared to state-of-the-art methods.
2 Related Work
Dataset with Motion Annotations. Most datasets [47, 54] focus on operation within the image space. However, camera and object motions are coupled in this space, while limiting the movement scope. On the other hand, only a few real datasets [51, 61, 57] provide camera pose, and these primarily focus on static scenes without dynamic objects. Some synthetic datasets [52, 3, 55, 46] provide pose annotations for both objects and camera, yet their 3D assets are predominantly human-centric, limiting category diversity. More discussions are in Sec 3.1.
Motion Control Methods. Most existing works [12, 19, 49, 59, 10] can only control either object motion or viewpoint change. For methods that support manipulation of both the camera and objects, Direct-a-Video [51] can only simulate basic movements and simultaneous control in MotionCtrl [47] often results in suboptimal outcomes as noted in its study [47]. More discussions are in Sec. 4.
3 SynFMC Dataset
3.1 Comparison with Existing Datasets
There is currently a lack of datasets that contain 6D poses of both objects and the camera [61, 57, 46]. As shown in Tab. 1, only a few real datasets [61, 57] provide estimated camera poses and are primarily limited to scenes without dynamic objects due to the suboptimal performance of estimation methods [31, 37, 60]. Some methods [54, 47] use in-the-wild videos with image-space object trajectory inferred by optical flow models [60], but this entangles object and camera motions while lacking 3D information like orientation. Synthetic datasets conveniently obtain pose information [3, 55, 46, 52], but most focus on human and are limited to small movements and simple camera motion patterns, limiting the ability to learn complex dynamic.
To facilitate model learning to control the motions in 3D-aware manner, it is essential to construct a new dataset with comprehensive object and camera poses. However, this is highly challenging in real world. First, capturing videos with complex, irregular object and camera motions is extremely difficult, typically requiring specialized equipment and expertise. Then, obtaining accurate pose estimation is difficult. Devices capable of capturing 6D poses for camera or objects are expensive and difficult to operate. Some studies [61, 47] attempt to obtain camera pose via estimation models. However, existing algorithms [60, 37] are time-intensive and often struggle with monocular videos containing dynamic objects. Besides, inferring 6D poses for general objects remains challenging. To address these limitations, we introduce SynFMC, a synthetic dataset generated using Unreal Engine [39], containing animations with diverse motion patterns and complete annotations.
$\bullet$ Difference from 360°360°360  °-Motion Dataset [10]. 1) We handle both static&dynamic cameras, enabling more complex shots than [10]’s static setup. 2) Our rule-based algorithm supports diverse (non-)horizontal object motions, unlike [10]’s GPT-derived horizontal-only ones. 3) Our environment & object assets extend beyond [10]’s terrestrial scenes.
$\bullet$ Difference from HumanVid-Syn Dataset [46]. 1) The object trajectories of [46] rely on predefined 3D motion assets (SMPL-X/skeleton), whereas our rule-based algorithm enables diverse patterns, e.g., in-place/(non-)horizontal motions in Fig. 32) [46] randomly samples camera positions within a semi-cylinder in front of human, while we achieve fine-grained control by decoupling camera motion (Fig. 4), supporting controllable and diverse movements. 3) [46] focuses on synthesizing single-human animation, whereas we support multi-object scenarios, generating richer dynamics.
3.2 Overview of SynFMC Dataset
SynFMC contains 62K videos divided into four groups: 15K static single-object, 15K static multi-object, 16K dynamic single-object, and 16K dynamic multi-object. Static means fixed object locations in world space while the camera remains movable. For diversity, the dataset includes common objects across various categories, such as humans, animals, plants and vehicles, as well as a wide range of environments like streets, grasslands, skies, oceans, etc. Additionally, SynFMC has diverse and complex multi-object and camera movements, covering not only basic motions but also shots that are challenging to achieve in real-world settings. Object assets are annotated with attributes such as speed and size by human annotators with the aid of MLLM [5] to ensure realistic motion simulation.
$\bullet$ Video Annotations. The proposed SynFMC dataset offers thorough annotations, including pose information for both objects and the camera, instance segmentation masks, depth maps, and detailed descriptions of content and motion, broadening its applicability across various fields [9, 8, 43].
Figure 3: Object motion types. The trajectory of stationary point is not presented in the figure, which is a fixed point in the space.
3.3 Data Generation Pipeline
$\bullet$ Asset Collection and Annotation. We collect diverse 3D assets, including environmens and objects. Environment assets span five types: groundnear groundskywater surface, and underwater. We collect high quality panoramic HDRI images from internet. For object assets, we select animated object assets from Objaverse-LVIS [7], Objaverse-XL [6], Mixamo [26], etc., covering diverse categories. Human annotators filter low-quality assets and verify/correct object properties (e.g., class, habitat, speed, size) queried by InternVL [5]. They also label motion type and provide description for each animation.
$\bullet$ Object Motion. To create realistic motions, we design trajectories based on Bézier curves in each motion segment, with rotations derived from tangent and normal vectors along the curve. Fig. 3 exemplifies motion types. Control points are constrained based on the object’s speed.
$\bullet$ Camera Motion. We decompose camera motion into 3 types: viewpointdistance, and height, see Fig. 4. (a) View-point controls camera’s orientation when capturing objects: front/back, left/right, and top. Start and end viewpoints of a motion segment are randomly assigned, with intermediate frames interpolated for smooth transitions, enabling camera’s motion range to encompass all orientations around the object. (b) Distance controls the horizontal distance between the camera and objects: zoom in/out and static. (c) Height controls the vertical distance: up/down and static. To maintain object visibility without centering it, the camera targets a randomly offset point near the object’s centroid.
1
3
2
4
5
7
6
8
9
Figure 4: Camera motion types. We decompose camera motion into 3 aspects. (a) Viewpoint controls camera orientation when capturing object. 1 - 5 present front/back, left/right and top perspectives. (b) Distance and (c) Height determine horizontal and vertical distance between camera and object, respectively. 6 - 9 are zoom in/out and up/down, respectively. The “static” types are omitted in (b) and (c), which stand for fixed distances.
Table 2: Comparison of FMC with other methods. FMC excels in controlling 6D poses of objects and camera with diverse motion patterns.
Methods Motion Condition Object Control Camera Control Dataset
AnimateDiff [11] Limited Patterns Dataset for Specific Motion Pattern
CameraCtrl [12] Camera Pose Diverse Patterns RealEstate10K [61]
VideoComposer [44] Image Space Trajectory Entanglement Entanglement LAION-400M [32] + WebVid [2]
DragNUWA [54] Image Space Trajectory Entanglement Entanglement WebVid+ VideoHD [54]
3DTrajMaster [10] Object Pose Limited Patterns 360°360°360  °-Motion [10]
Direct-a-Video [51] Image Space Trajectory + Camera Type Entanglement Limited Patterns MovieShot [29]
MotionCtrl [47] Image Space Trajectory + Camera Pose Entanglement Diverse Patterns RealEstate10K + WebVid
FMC (ours) Object Pose + Camera Pose Diverse Patterns Diverse Patterns SynFMC (ours)
Figure 5: Motion trajectories. The examples of horizontal and non-horizontal motions indicate the variety and complexity of the trajectories generated by our rule-based algorithm.
$\bullet$ Generation of Multi-Object Scenes. Objects from the same environment with comparable sizes and speeds are selected. The first object’s trajectory is created using the way described earlier, while subsequent trajectories are derived based on the preceding object’s path with a reasonable offset. For camera motion, we randomly track an object in each motion segment, with the trajectory generated using the same method as previously outlined.
$\bullet$ Rendering. This stage creates synthetic videos based on selected assets and motion types of objects and the camera. We divide a complete motion trajectory into many small segments. For each segment, we randomly select animations for objects and generate trajectories according to their annotated motion types. Similarly, random combinations of camera motion types are applied across different segments, allowing the generated videos to encompass diverse motion patterns as in real-world scenarios. These segments are seamlessly combined to form a complex and diverse global trajectory, as shown in Fig. 5. Finally, the selected assets and poses of the camera and objects are imported into Unreal Engine [39] to get 3D animation sequence for rendering. Fig. 1 shows an example trajectory segment.
4 The Proposed Approach
Given $N$-length camera poses $\mathcal{C}_{RT}$ = ${RT}_{cam}^{1:N}$, object poses $\mathcal{O}_{RT}$ = $\{RT_{obj_{i}}^{1:N}\}_{i=1}^{N_{o}}$ of $N_{o}$ objects in the global coordinate system, and content description $\mathbf{C}_{p}$, we aim to generate the video that reveals correct motion in real world.
As shown in Tab. 2, most methods cannot independently or jointly control object and camera movements in a 3D-aware manner. For example, AnimateDiff [11] and CameraCtrl [12] only support camera control. Methods [44, 54, 51, 47] using image-space trajectories faces motion entanglement issues and can’t control the orientation. Although Direct-a-Video [51] introduces several camera types and allows explicit control, it is limited to simple motion patterns. MotionCtrl [47], the closest to ours, trains two motion modules separately but lacks comprehensive 6D pose annotations, resulting in suboptimal simultaneous control of camera and object motions. Additionally, it simply applies standard diffusion loss [13] when training motion modules, which further hinders its ability to disentangle camera and object motions within a video. FMC trained on SynFMC introduces Camera Motion Controller (CMC) and Object Motion Controller (OMC) to address these limitations. OMC receives 6D pose and coarse mask of the object to perceive its spatial location and orientation, achieving a realistic appearance from various viewpoints. The training objectives enable FMC to disentangle the motion effects of objects and the camera in the video, allowing independent or joint 3D-aware control of camera and object motions.
$\bullet$ Preliminary. 1) T2V diffusion models [11, 14, 58, 56, 36, 15] add Gaussian noise $\epsilon$ to image sequence $\mathbf{z}_{0}^{1:N}$ in training, resulting in noisy latents $\mathbf{z}_{t}^{1:N}$ at $t$ time step. Network $\varepsilon_{\theta}$ then is trained to infer the injected noise from current latents. 2) LoRAs [17, 11] are used to learn different content styles. We apply this to bridge the synthetic-real video domain gap. 3) Following [12], we use plücker embedding to represent camera pose for geometric interpretation.
Figure 6: Domain LoRA. We sample the first frame of generated videos under without and with Domain LoRA settings.
Figure 7: The architecture of FMC. In the first stage, we randomly sample the images from synthetic videos and update the parameters from injected Domain LoRA. Next, the modules from CMC are learned. It consists of two parts: Camera Encoder and Camera Adapter, where the Camera Adapter is introduced into the temporal modules. Finally, we train the Object Encoder from OMC. It receives the 6D object pose features, which are repeated in the corresponding object region. We use Gaussian blur kernel centered at the centroid to prevent the need of precise masks. Then, the output is multiplied by the coarse masks to modulate the features in the main branch.
4.1 Free-Form Motion Control
Fig. 7 shows the overall architecture of the proposed FMC method. We train it in 3 stages. First, Domain LoRA [17] are injected into spatial blocks to adapt to rendered content, with temporal modules inactive and images randomly sampled from synthetic data. Fig. 6 shows the effectiveness of this stage in bridging the domain gap. Then, CMC is trained to learn camera motion, introducing temporal modules and loading LoRA from the previous step. Finally, OMC is trained to decouple object dynamics from camera motion, with other parameters frozen. During inference, LoRA modules are dropped to maintain the quality of base model.
$\bullet$ Camera Motion Controller (CMC). It consists of two parts as shown in Fig. 7. Camera Encoder receives plücker embeddings, where the initial camera pose (translation values are set to 0) and the relative camera poses are used for the first and subsequent frames respectively. The initial pose helps to determine the perspective at the start time. Then, the outputs are processed by Camera Adapter to modulate the features in temporal blocks. Due to the dynamic of the background being only affected by camera motion, camera loss $L_{cam}$ is applied in this stage:
[…]
where $\theta_{c}$ is the parameters of CMC. $\mathcal{M}_{bg}$ is background mask and $\lambda_{c}$ is the weighting factor. $L_{cam}$ makes camera motion more accurate by concentrating on the background.
$\bullet$ Object Motion Controller (OMC). The Object Encoder of OMC receives 6D object pose information to adjust the features in spatial modules from several downsample blocks [30, 47]. Specifically, the poses relative to the camera in each frame are duplicated within the respective object region while the others are set to 0. Then, the pose features concatenated with the foreground mask are fed to OMC. In this manner, the poses from different objects can be aggregated in a single input. Besides, we leverage the Gaussian blur kernel centered at the object centroid to avoid users offering precise masks. Then, the outputs from OMC are multiplied by the coarse masks and added to the spatial features in the main branch, preventing to impair the background content. During inference, the size of the kernel can be approximated based on the object’s size (specified by the user) and its distance from the camera. object loss $L_{obj}$ is applied to make OMC focus on the object region:
[…]
where $\theta_{o}$ indicates the parameters from OMC. $\mathcal{M}_{fg}$ is foreground mask and $\lambda_{o}$ is the weighting factor. $L_{obj}$ improves the appearance quality of dynamic objects.
5 Experiments
$\bullet$ Implementation Details. The proposed FMC is based on AnimateDiff V3 [11], trained with 16-length $256\times 384$ videos, Adam optimizer with a learning rate of $1e^{-4}$. Domain Adapter is trained with 8K iterations in a batch size of 128. CMC and OMC are trained with 50K iterations with a batch size of 8. $\lambda_{c}$ and $\lambda_{o}$ are set to 0.6 and 0.3. In experiments, we find that 26K data samples suffice for our method to learn 3D-aware motion control.
$\bullet$ Evaluation Metrics. Following [11, 47], we use FID [33] to evaluate visual quality, FVD [40] for temporal coherence, and CLIPSIM [28] to measure semantic similarity with text. For camera motion, we follow [12] to use CamTransErr and CamRotErr. For object motion, ObjTransErr and ObjRotErr are introduced. We first use depth estimation model [50] to obtain the depth at object centroid and determine its global position based on camera pose, then fit a trajectory curve to find tangent and normal vectors at each time step, allowing for rotation derivation. Given scale information, we apply appropriate scaling to the translation error calculation.
Figure 8: Independent controls over camera and object motions. Results in (a) reveal that all methods [12, 47] effectively reflect the camera conditions. For object motion, the compared methods [47, 51] fail to maintain a stationary camera as shown in green boxes from (b) (e.g., movement of flower in row 5). Furthermore, they also present low fidelity of object orientation (3D axes in conditions).
5.1 Comparisons with State-of-the-Art Methods
We first compare independent controls over camera motion and object motion with previous methods [47, 12]. Then, we demonstrate FMC ’s superior performance in simultaneous control. Finally, we showcase additional examples across different scenes to validate the effectiveness of SynFMC and FMC. For fairness, we compare with U-Net based methods. More examples are in supplementary.
$\bullet$ Independent Control of Camera Motion. MotionCtrl [47] and CameraCtrl [12] are selected for this comparison as they accept explicit camera information. In Fig. 8(a), we simulate two camera motions and scale the translation to fit the input range required by these methods. FMC and the compared methods effectively reflect the input conditions. The CamTransErr & CamRotErr in Tab. 3 also show that FMC achieves comparable results in controlling camera.
Table 3: Quantitative comparison of our proposed method FMC with AnimateDiff [11], CameraCtrl [12], and MotionCtrl [47].
Method AnimateDiff CameraCtrl MotionCtrl FMC (ours)
FID $\downarrow$ 149.61 137.96 125.52 133.42
FVD $\downarrow$ 868.97 805.25 952.31 846.51
CLIPSIM $\uparrow$   29.33    29.21   26.83   31.01
CamTransErr $\downarrow$ -    18.16    17.84   18.12
CamRotErr $\downarrow$ -      0.94     1.11     1.03
ObjTransErr $\downarrow$ - -    80.66   42.25
ObjRotErr $\downarrow$ - -      1.77     0.96
Figure 9: Simultaneous control over camera and object motions. MotionCtrl [47] struggles to generate realistic object dynamics, causing objects to disappear from view, whereas our FMC achieves high-quality simultaneous control.
$\bullet$ Independent Control of Object Motion. For object control, we compare with MotionCtrl [47] and Direct-a-Video [51]. For these methods, we project the global trajectory into image space using camera and object pose information. As shown in Fig. 8(b), the compared methods fail to maintain a stationary camera (showing dynamic movement in the background), indicating that image space trajectories entangle the object and camera motions. For example, the 2nd example of Fig. 8(b), Direct-a-Video [51] shows the change of flower location caused by dynamic camera. Our method effectively alleviates this issue by incorporating static camera poses as constraints. Furthermore, since OMC receives 6D pose of objects, FMC can achieve high fidelity of object orientation with input condition.
Table 4: User study in quality, text similarity, and motion fidelity.
Method CameraCtrl [12] MotionCtrl [47] FMC (ours)
Quality Score 0.88 0.89 0.91
Text Similarity Score 0.84 0.81 0.95
Camera Motion Score 0.95 0.93 0.95
Object Motion Score - 0.53 0.98
Table 5: Quantitative results in ablation study.
Metrics CamTransErr CamRotErr ObjTransErr ObjRotErr
MotionCtrl (w/o $\mathcal{C}_{RT}$) 18.24 1.08 78.82 1.65
MotionCtrl (w/ $\mathcal{C}_{RT}$) 18.24 1.08 55.33 1.26
FMC (w/o $L_{cam}$) 20.35 1.19 - -
FMC (w/o $L_{obj}$) 18.12 1.03 46.62 1.15
FMC 18.12 1.03 42.25 0.96
$\bullet$ Simultaneous Control of Camera and Object Motions. We explore the combination of both camera and object control signals. Since Direct-a-Video [51] supports only basic camera motion, we choose MotionCtrl [47] as the comparison method. We randomly simulate movements for both the object and the camera, resulting in varied trajectories. As shown in Fig. 9, videos generated by FMC more faithfully align with the specified conditions. While MotionCtrl captures the camera’s motion, it struggles to generate realistic object dynamics. These results demonstrate the effectiveness of FMC in achieving simultaneous control of camera and object motions. The object error metrics in Tab. 3 show that our method achieves better results in object motion control. Additionally, FMC achieves higher scores in user study as shown in Tab. 4, outperforming previous methods in quality and motion fidelity.
Fig. 10 shows video generation results across 4 different cases: static single-object, dynamic single-object, static multi-object, and dynamic multi-object. Thanks to the diversity of motion patterns in SynFMC, FMC effectively learns a range of diverse, advanced, and complex shots. In the 2nd row of Fig. 10, for example, the camera initially captures the person from the front and then follows from behind. The last two rows demonstrate the performance in multi-object scenarios, where the relative motion between objects and the camera conforms closely to the input conditions.
5.2 Ablation Study
Figure 10: Simultaneous control results of FMC in different cases. The complicated case in row 2 shows that our method learns complex shot, where the camera first captures a skier from the front and then follows him from behind.
Figure 11: Simultaneous control results of MotionCtrl [47] trained on SynFMC without and with camera pose during training.
$\bullet$ SynFMC Dataset. To validate the effectiveness of complete camera and object pose annotations in SynFMC and evaluate the dataset generalization, we train MotionCtrl [47] on it, adapting the annotations to its input format. We first train the camera module, then optimize the object module in two ways: without camera poses, as in [47], and with camera poses, as our FMC. As shown in Fig. 11, incorporating known camera poses when optimizing object module allows the object to follow the input trajectory more accurately, reducing the risk of it leaving the field of view. Object motion errors in Tab. 5 further highlight the benefits of using complete camera and object pose annotations.
$\bullet$ OMC. As shown in the first 2 rows of Fig. 12 and object motion errors in Tab. 5, FMC outperforms MotionCtrl [47] trained on SynFMC in motion accuracy. This improvement is brought by OMC’s ability to process 6D poses, generating more realistic object appearances based on orientation and object’s distance from the camera, reflected by the size of coarse masks. [47]’s object motion control module can only handle 2D image-space trajectories without pose and distance information, limiting alignment accuracy with input.
Figure 12: Results of different settings in the ablation study. The first row is MotionCtrl [47] trained on SynFMC.
$\bullet$ Training Objectives. We conduct two experiments to assess the impact of $L_{cam}$ in Eq. 1 and $L_{obj}$ in LABEL:eq:obj-ddpm-loss. First, we train CMC using only the standard diffusion loss [13, 45]. As shown in the 3rd row of Fig. 12, model without $L_{cam}$ tends to shift foreground objects to achieve similar relative motion, which does not accurately match the input pose. The camera motion error in Tab. 5 underscores the effectiveness of $L_{cam}$. Besides, training OMC without $L_{obj}$ leads to undesired object appearances, as shown in the 5th row of Fig. 12, with object motion error in Tab. 5 confirming the benefit of $L_{obj}$.
$\bullet$ Adaptability with Different T2I Personalized models. As shown in Fig. 2, FMC can be adapted to various personalized backbones [38, 41, 30], showing that our proposed dataset SynFMC and corresponding training strategy do not impair the model’s original generative capabilities.
6 Conclusion
This work introduces SynFMC, a dataset with comprehensive 6D pose information and diverse assets, offering both standard and complex shots that are difficult to capture in real life, with trajectories resembling real-world scenarios. With SynFMC, the proposed method FMC enables independent or joint 3D-aware control of object and camera motions within a single video. Experimental results demonstrate the effectiveness of both SynFMC dataset and FMC.
Limitations. Our method’s ability to control complex motions of multiple objects remains limited. Better metrics are needed to more accurately evaluate object motion. In the future, additional input modalities, e.g., images, are desired to customize motion videos for reference subjects.
Acknowledgement.
This project was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 62472104. This work was supported by Damo Academy through Damo Academy Innovative Research Program. Dr Tao’s research is partially supported by NTU RSR and Start Up Grants. This work was in part supported by National Natural Science Foundation of China (Grant No. 62276067).